Calculate the number of.pounds of CO2 released into atmosphere when a 20.0 gallon tank of gas is burned in a car engine. Assume that gas Is octane C8H18 and the density of gas is 0.692g.mL-1. what is CO2 released in pounds ?


1
Expert's answer
2020-11-09T14:08:17-0500

20 gallons = 75.7 L = 75700 mL

m(C8H18) = 75700*0.692 = 52384 g

n(C8H18) = 52384/(12*8+18) = 460 mol

n(CO2) = 460*8 = 3680 mol

m(CO2) = 3680*(12+16*2) = 161920 g = 357 pounds

Answer: 357 pounds